3.34 A closed cycle consists of the three processes listed in the table below. Fill in the missing values if all the quantities are in Btu/Ib (Hint: A closed cycle returns to its starting point.) Process q Au 200100 100 2+3 3- y150 300 Simple & clear please
Answer #1

we know that by first law of thermodynamics

\Delta Q = \Delta W + \Delta U

and for a cyclic condtion

\int dW = \int dQ

using above equations, the solution can be found easily

which is given in figure below

200 = too + 2 3-2 y=150 2

hit like if answer helped.